OpenStudy (mertsj):

If g(x)+xsin(g(x))= x^2 find g'(0)

OpenStudy (jamesj):

Differentiating this expression: g(x)+xsin(g(x))= x^2 g'(x) + sin(g(x)) + x.g'(x) cos(g(x)) = 2x Now set x = 0 and we have g'(0) + sin(g(0)) + 0 = 0 hence g'(0) = - sin(g(0))